Grade 11Mechanics

From a point on ground at a distance 2 meters from the foot of a vertical wall a ball is thrown at an angle of 45degree which just clears the top of the wall.and afterward strikes the ground at a distance 4m on the other side.The height of wall is

Anirudh Chourey
7 Years agoGrade 11

1 Answer

Khimraj
7 Years ago
Total range = 2 + 4 = 6
Horizontal velocity = range/ time of take off
U cos45 degree = 6 x g / 2 U sin45 degree
U= 8 m/s
Time= distance/ velocity
6/ 8 cos 45 degree 
= 0.75\sqrt{2} sec
Displacement= h= U sintheta x t  - 1/2 gt²
h = 8 sin 45*0.75\sqrt{2}- 4.9*2*0.75^2 
h = 2/3 m/s